# Schema-Driven Information Extraction from Heterogeneous Tables

### v3

@@ Abstract
In this paper, we explore the question of whether large language models can support cost-efficient information extraction from tables. We introduce schema-driven information extraction, a new task that transforms tabular data into structured records following a human-authored schema. To assess various LLM’s capabilities on this task, we present a benchmark comprised of tables from four diverse domains: machine learning papers, chemistry literature, material science journals, and webpages. We use this collection of annotated tables to evaluate the ability of open-source and API-based language models to extract information from tables covering diverse domains and data formats. Our experiments demonstrate that surprisingly competitive performance can be achieved without requiring task-specific pipelines or labels, achieving F1 scores ranging from 74.2 to 96.1, while maintaining cost efficiency. Moreover, through detailed ablation studies and analyses, we investigate the factors contributing to model success and validate the practicality of distilling compact models to reduce API reliance. [^1] Code and data are available at our GitHub repository.
@@ 1 Introduction
Vast quantities of data are locked away in tables found in scientific literature, webpages, and more. These tables are primarily designed for visual presentation, and the underlying data is typically not available in any structured format, such as a relational or graph database. Some table collections have simple or uniform structures Cafarella et al. (2008), making them easy to convert to relational data, for example, Wikipedia tables Lebret et al. (2016); Iyyer et al. (2017), however a lot of information is stored in tables with complex and varied layouts, such as tables of experimental data presented in scientific literature.
@@
Prior work on extracting structured data from tables has focused on developing custom pipelines for each new table format or domain, for example extracting machine learning leaderboards from  result tables Kardas et al. (2020). Importantly, the development of these specialized pipelines necessitates domain-specific labeled data, which not only incurs a significant cost in collection for every new extraction task but also constrains their applicability outside the originating domain.
@@
In this paper, we demonstrate that large language models can enable accurate domain-independent extraction of data from heterogeneous tables. To show this, we present a new formulation of the table extraction problem, which we refer to as Schema-Driven Information Extraction. In Schema-Driven IE, the only human supervision provided is a schema that describes the data model, including the target attributes and their data types, formulated in a JSON format. [^2] JSON is chosen as the output format for two main reasons: 1) its widespread use ensures a significant representation in the LLM’s pre-training corpus, which is crucial for optimizing model performance; and 2) its simplicity in parsing and processing, especially its support for one-line output, makes it advantageous for outputs spanning multiple cells, offering a clear benefit over indent-based formats like YAML. Given an extraction schema, and a table as input, the model then outputs a sequence of JSON objects, each of which describes a single cell in the table and adheres to the user-provided schema. For example, as demonstrated in Figure 1, a domain expert outlines the attributes of interest related to experimental result cells in a machine learning table, and the model extracts JSON objects following this schema.
@@
To evaluate the ability of LLMs to perform Schema-Driven IE, we introduce a new benchmark consisting of table extraction datasets in four diverse domains: machine learning papers, chemistry literature, material science journals, and webpages - each of which has a different data format (, XML, CSV, and HTML, respectively). We curate and annotate data for the first two domains, while adapting existing datasets for the latter two.
@@
We then use this benchmark to analyze the performance of open-source and proprietary LLMs. We find that proprietary models perform well across diverse domains and data formats. For example, GPT-4 OpenAI (2023) and code-davinci Chen et al. (2021), are capable of accurate table extraction (ranging from 74.2 to 96.1 F1), given only a relevant data schema as supervision. This performance is comparable to fully supervised models, which operate at an F1 range of about 64.1 to 96.1. We also present a number of analyses on various factors that are key to achieving good performance while minimizing inference costs, including retrieving text from outside the table, in addition to an iterative error recovery strategy. Moreover, we demonstrate the utility of Schema-Driven IE by evaluating performance on the downstream task of leaderboard extraction from machine learning papers (Kardas et al., 2020).
While open-source models have yet to match the performance of their proprietary counterparts, our analysis reveals that recent models like CodeLlama-instruct-13B (Rozière et al., 2023) show significant progress, e.g., on ML tables, its performance is comparable to that of GPT-3.5-turbo.
Furthermore, we conduct comprehensive ablation studies and analyses to explore model performance nuances, and demonstrate the feasibility of distilling efficient table extraction models without compromising performance. By introducing a new benchmark and a baseline for Schema-Driven IE, our goal is to encourage the creation of future open-source models capable of executing this task independently of proprietary APIs.
@@ 2 Schema-Driven Information Extraction
We present a new task for extracting structured records from tables. As shown in Figure 1, the task input contains two elements: 1) a table with numerous cells, optionally supplemented with contextual text, e.g., retrieved paragraphs from the same document; and 2) an extraction schema that outlines target attributes and their data types for various record types (implemented as JSON templates). Given the input, the model generates a sequence of JSON objects, where each object corresponds to a cell in the table and contains key-value pairs for the pre-defined attributes of a specific record type.
@@
Despite explicit instructions, we found that models often fail to generate JSON records for all the cells in a single inference pass. Instead, models deviate from the instructed cell traversal order, leading to partial extraction of the input table’s cells. To mitigate this, we use an iterative error recovery strategy. As shown on the right side of Figure 2, we detect deviations from the instructed left-right, top-down order by comparing predicted cell values with those from a rule-based cell detector. Then, we truncate the LLM’s output to the point of deviation, and re-prompt the model with the truncated sequence, adding the value of the next target cell. This strategy guides the model to adhere to the instructed order, and continues iteratively until all records are generated. In Section 4.4, we show that this approach achieves comparable performance to cell-by-cell prompting, while significantly reducing inference costs. For more details on InstrucTE, including prompt formulation and cell detectors, please refer to Appendix A.
@@ 3 The Schema-to-Json Benchmark
We now present the details of our benchmark, Schema-to-Json, which is designed to assess the capabilities of LLMs to extract data from tables, adhering to a predefined schema. This benchmark contains tables from four diverse domains: machine learning papers, chemistry literature, material science journals, and webpages. Each domain features a unique textual format, namely, , XML, CSV, and HTML, requiring models to generalize across domains and formats. For ML tables, we add relevant paragraphs from the same documents to provide additional context, testing the models’ capacity to jointly understand tabular and textual data. We manually annotate datasets for the first two domains and adapt pre-existing datasets into our unified format for the latter two. Statistics of the four datasets are summarized in Table 1.
@@
We create a manually annotated dataset focused on tables from arXiv ML papers, emphasizing numeric cells that are classified into four categories: Experimental Results, Hyper-parameters, Data Statistics, or Other. Extraction attributes are pre-defined for the first three categories; for instance, Result records incorporate textual attributes such as evaluation metric (e.g., F1) and dataset (e.g., SQuAD), as shown in Figure 1. We collect papers from three subfields: Machine Learning, Computer Vision, and Natural Language Processing, and randomly select five tables from each paper (including those in appendices) for budget reasons. [^3] We limit the paper selection to those published between Oct. and Nov. 2022 to avoid contamination with GPT-4 (0613) and GPT-3.5, which were trained on data until Sep. 2021. We employ computer scientists with ML backgrounds for annotation, and evaluate inter-annotator agreement (IAA) score by calculating Table-F1 (a metric detailed in Section 4.1) on double-annotated tables, treating one set of annotations as gold labels and the other as predictions. This method yields a Table-F1 score of 96.6 when applying thresholded token-level F1 for attribute matching. For additional information on the ML tables, including predefined attributes and the annotation process, please refer to Appendix B.
PubMed Chemistry Tables
We also annotate a new dataset of PubMed tables describing the physical properties of chemical compounds. The automated extraction of physical properties from such tables could provide substantial real-world benefits, for example collecting much-needed data for training ML models that can support inverse molecular design (Kim et al., 2018) and thus accelerating the drug design process (Fields, 2019; Stokes et al., 2020). Here, we focus on cells concerning five important physical properties identified by chemists: $\text{IC}_{50}$, $\text{EC}_{50}$, $\text{GI}_{50}$, $\text{CC}_{50}$, and MIC. [^4] https://www.sciencedirect.com/topics/pharmacology-toxicology-and-pharmaceutical-science/ic50 Three common attributes are manually extracted from tables for all properties: unit, treatment (experimental compound), and target (measured biological entity, e.g., a gene expression). Similar to the ML tables, domain experts annotate JSON records for relevant cells, and Table-F1 calculated on double-annotated tables is used as the IAA score. A Table-F1 score of 91.0 is achieved when applying thresholded token-level F1 for attribute matching, underscoring the reliability of the dataset.
DisCoMat (Gupta et al., 2022)
We incorporate DisCoMat, an existing dataset focusing on glass composition tables from Elsevier material science journals. The task is to extract tuples comprising (material, constituent, percentage, unit) from given tables. We adapt DisCoMat to fit our Schema-Driven IE framework by grounding the percentage element to numeric cells in the table and considering the other elements as attributes. The model is tasked to identify numeric cells representing constituent percentages and predict the associated three attributes. We refer readers to Gupta et al. (2022) for more details of DisCoMat. [^5] In the released corpus, tables are represented as matrices; we, therefore, transform them into CSV tables (using the pipe symbol ”—” as the delimiter) prior to feeding them into LLMs.
SWDE (Hao et al., 2011)
@@
# papers (web.) 25 16 2,536 80
# tables (pages) 122 26 5,883 1,600
# anno. records 3,792 1,498 58,481 1,600
# records / table 31.1 57.6 9.9 1
@@ 4 Experiments
We evaluate the capability of various LLMs to perform schema-driven information extraction, in addition to full fine-tuning using our benchmark. For ML and chemistry tables, we use a subset of 10 and 7 randomly sampled papers separately for model development, which facilitates the training of supervised models, thereby enabling comparison with a schema-driven approach. For the two pre-existing datasets, we follow the data splits used in the original experiments.
4.1 Evaluation Metrics
We introduce Table-F1, a new reference-based evaluation metric gauging attribute prediction performance within a table. Table-F1 represents the harmonic mean of precision and recall, with precision being the ratio of correctly predicted attributes to total predicted attributes. At the attribute level, we consider two metrics: token-level F1 and exact match (EM). For token-level F1, a prediction is deemed correct if the score exceeds a specific threshold, which is determined by maximizing the alignment between model predictions and human judgments on the dev set (see Appendix C). An example of Table-F1 calculation is shown in Figure 7 in the appendix. We report macro-averaged Table-F1 given the wide variance in table sizes.
For DisCoMat and SWDE, we use the metrics specified in the original papers to support comparisons with prior work. In the case of DisCoMat, we report Tuple-F1 (Gupta et al., 2022), where a predicted 4-element tuple is considered correct only if it exactly matches the gold tuple. For SWDE, we report Page-F1 (Hao et al., 2011), which measures the number of pages where the attributes are accurately predicted. [^6] Notably, SWDE primarily focuses on identifying textual HTML nodes containing attribute values rather than exact text spans, so we use token-level F1 to identify the most relevant HTML node for each extracted attribute.
@@ 4.2 Baselines & Implementation Details
We evaluate the capability of multiple LLMs to perform Schema-Driven IE, including API-based GPT-4 and GPT-3.5 models and open-source models, such as Llama2-Chat-13B (Touvron et al., 2023b), CodeLlama-instruct-13B (Rozière et al., 2023), StarCoder-15.5B (Li et al., 2023), LLaMA-7B (Touvron et al., 2023a), and Alpaca-7B (Taori et al., 2023). We also frame Schema-Driven IE as a TableQA problem, applying multi-choice and extractive QA prompts for template selection and cell attribute prediction, respectively. Furthermore, we also evaluate T5-11B (Raffel et al., 2020) and TaPas (Herzig et al., 2020), a table-specialized LM. For implementation details of InstrucTE and other methods, see Appendix D.
@@ 4.3 Main Results
Despite a noticeable gap when compared to API-based LLMs, open-source models, like CodeLlama-instruct-13B, show promising results in ML and web domains, achieving 60.0 Table-F1 and 91.7 Page-F1 on ML tables and SWDE, respectively. This success is likely due to these domains being more represented in the model’s pre-training corpus, suggesting that enhancing the pre-training data in less represented domains, such as chemistry and material science, may be an avenue for narrowing the gap with API-based models.
@@ 4.4 Ablation Studies
Prompt Components & Error Recoverery
@@ 4.5 Performance Analysis
To gain more insights about InstrucTE’s performance, we now discuss a set of key questions.
What types of tables can be better processed by InstrucTE?
We analyze InstrucTE’s performance against table complexity on ML tables, considering two factors: 1) the total cell count, and 2) the number of multi-row/column cells, gauging structural complexity. The Pearson correlation coefficients with InstrucTE’s performance (using GPT-4) are 0.096 and 0.128, respectively, suggesting the model’s effective processing across various table sizes and complexities. Qualitative analysis reveals that InstrucTE performs best with tables containing common terminologies and standard content, particularly when contextually supported by the surrounding text. For example, it accurately identifies "Result" cells featuring well-known datasets (e.g., SQuAD) and metrics (e.g., F1 score). Yet, the challenge lies in quantifying the familiarity of terminology and content, given the model’s black-box nature, and addressing it represents a compelling direction for future research.

### v4

@@ Abstract
In this paper, we explore the question of whether large language models can support cost-efficient information extraction from tables. We introduce schema-driven information extraction, a new task that transforms tabular data into structured records following a human-authored schema. To assess various LLM’s capabilities on this task, we present a benchmark comprised of tables from four diverse domains: machine learning papers, chemistry literature, material science journals, and webpages. We use this collection of annotated tables to evaluate the ability of open-source and API-based language models to extract information from tables covering diverse domains and data formats. Our experiments demonstrate that surprisingly competitive performance can be achieved without requiring task-specific pipelines or labels, achieving F1 scores ranging from 74.2 to 96.1, while maintaining cost efficiency. Moreover, through detailed ablation studies and analyses, we investigate the factors contributing to model success and validate the practicality of distilling compact models to reduce API reliance. [^1] Our code and data are available at https://github.com/bflashcp3f/schema-to-json.
@@ 1 Introduction
Vast quantities of experimental data are locked away in tables found in scientific literature. These tables are primarily designed for visual presentation, and the underlying data is typically not available in any structured format, such as a relational or graph database. Some table collections have simple or uniform structures Cafarella et al. (2008), making them easy to convert to relational data, for example, Wikipedia tables Lebret et al. (2016); Iyyer et al. (2017), however a lot of information is stored in tables with complex and varied layouts, such as tables of results in papers found on arXiv.org.
@@
Prior work on extracting data from tables has focused on developing custom pipelines for each new table format or domain, for example extracting machine learning leaderboards from  result tables Kardas et al. (2020). Importantly, the development of these specialized pipelines necessitates domain-specific labeled data, which not only incurs a significant cost in collection for every new extraction task but also constrains their applicability outside the originating domain.
@@
In this paper, we show how LLMs can enable accurate domain-independent extraction of data from heterogeneous tables. We present a new formulation of the table extraction problem, which we refer to as Schema-Driven Information Extraction. In Schema-Driven IE, the only human supervision provided is a schema that describes the data model, including the target attributes and their data types, formulated in a JSON format. [^2] JSON is chosen as the output format for two main reasons: 1) its widespread use ensures a significant representation in the LLM’s pre-training corpus, which is crucial for optimizing model performance; and 2) its simplicity in parsing and processing, especially its support for one-line output, makes it advantageous for outputs spanning multiple cells, offering a clear benefit over indent-based formats like YAML. Given an extraction schema, and a table as input, the model then outputs a sequence of JSON objects, each of which describes a table cell. For example, as demonstrated in Figure 1, a domain expert outlines the attributes of interest related to result cells in a machine learning table, and the model extracts JSON objects following this schema.
@@
To evaluate the ability of LLMs to perform Schema-Driven IE, we introduce a new benchmark consisting of table extraction datasets in four diverse domains: machine learning papers, chemistry literature, material science journals, and webpages - each of which has a different data format (, XML, CSV, and HTML, respectively). We curate and annotate new datasets for the first two domains, while adapting existing datasets for the latter two.
@@
Using this newly developed benchmark, we analyze the performance of open-source and proprietary LLMs. We find that state-of-the-art proprietary models are capable of accurately extracting data from diverse domains and table formats without supervision. For example, GPT-4 OpenAI (2023) and code-davinci Chen et al. (2021), are capable of accurate table extraction (ranging from 74.2 to 96.1 F1), given only a relevant data schema as input to define the task. This performance is comparable to fully supervised models, which operate at an F1 range of about 64.1 to 96.1. We also present a number of analyses on various factors that are key to achieving good performance while minimizing inference costs, including retrieving text from outside the table, in addition to an iterative error recovery strategy. Moreover, we demonstrate the utility of Schema-Driven IE by evaluating performance on the downstream task of leaderboard extraction from machine learning papers (Kardas et al., 2020).
@@ 2 Schema-Driven Information Extraction
We now describe Schema-Driven IE, a new task that extracts structured records from tables with minimal supervision. As shown in Figure 1, the task input contains two elements: 1) a table with numerous cells, optionally supplemented with contextual text, e.g., retrieved paragraphs from the same document; and 2) an extraction schema that outlines target attributes and their data types for various record types (implemented as JSON templates). Given the input, the model generates a sequence of JSON objects, where each object corresponds to a cell in the table and contains key-value pairs for the pre-defined attributes of a specific record type.
@@
Despite explicit instructions, we found that models often fail to generate JSON records for all the cells in a single inference pass. Instead, models often deviate from the instructed cell traversal order, leading to partial extraction of the input table’s cells. To mitigate this, we use an iterative error recovery strategy. As shown on the right side of Figure 2, we detect deviations from the instructed left-right, top-down order by comparing predicted cell values with those from a rule-based cell detector. Then, we truncate the LLM’s output to the point of deviation, and re-prompt the model with the truncated sequence, adding the value of the next target cell. This process is repeated until all records are generated. Using identified cells as a scaffold, this strategy helps the model adhere to the instructed order, significantly improving performance despite potential propagated errors in cell identification. In Section 4.4, we show that our approach is much more cost-efficient than cell-by-cell prompting while achieving similar performance. For more details on InstrucTE, including prompt formulation and cell detectors, please refer to Appendix A.
@@ 3 The Schema-to-Json Benchmark
We now present the details of our benchmark, Schema-to-Json, which is designed to assess the capabilities of LLMs to extract data from tables, adhering to a predefined schema. This benchmark contains tables from four domains: machine learning papers, chemistry literature, materials science journals, and webpages. Each domain features a unique textual format, namely, , XML, CSV, and HTML, requiring models to generalize across domains and formats. For ML tables, we add relevant paragraphs from the same documents to provide additional context, testing the models’ capacity to jointly understand tabular and textual data. We manually annotate datasets for the first two domains and adapt pre-existing datasets into our unified format for the latter two. Statistics of the four datasets are summarized in Table 1.
@@
We create a manually annotated dataset focused on tables from arXiv ML papers, emphasizing numeric cells that are classified into four categories: Results, Hyper-parameters, Data Statistics, or Other. Extraction attributes are pre-defined for the first three categories; for instance, result records incorporate textual attributes such as evaluation metric (e.g., F1) and dataset (e.g., SQuAD), as shown in Figure 1. To avoid data contamination with top models like GPT-4 (0613), [^3] According to OpenAI website, GPT-4 (0613) was trained on data until Sep. 2021. we collected papers published after the knowledge cutoff (between October and November 2022) from three subfields: Machine Learning, Computer Vision, and Natural Language Processing. Five tables were randomly selected from each paper, including appendices. We employ computer scientists with ML backgrounds for annotation, and evaluate inter-annotator agreement (IAA) score by calculating F1 (see Section 4.1 for details) on double-annotated tables, treating one set of annotations as gold labels and the other as predictions. This method yields an F1 score of 96.6 when applying thresholded token-level F1 for attribute matching. For additional information on ML tables, including predefined attributes and the annotation process, please refer to Appendix B.
@@
# papers (web.) 25 16 656 80
# tables (pages) 122 26 1,031 1,600
# anno. records 3,792 1,498 9,036 1,600
# records / table 31.1 57.6 8.8 1
@@
PubMed Chemistry Tables
We also annotate a new dataset of PubMed tables describing the physical properties of chemical compounds. The automated extraction of physical properties from such tables could provide substantial real-world benefits, for example collecting much-needed data for training ML models that can support inverse molecular design (Kim et al., 2018) and thus accelerating the drug design process (Fields, 2019; Stokes et al., 2020). Here, we focus on cells concerning five important physical properties identified by chemists: $\text{IC}_{50}$, $\text{EC}_{50}$, $\text{GI}_{50}$, $\text{CC}_{50}$, and MIC. [^4] https://www.sciencedirect.com/topics/pharmacology-toxicology-and-pharmaceutical-science/ic50 Three common attributes are manually extracted from tables for all properties: unit, treatment (experimental compound), and target (measured biological entity, e.g., a gene expression). Similar to the ML tables, domain experts annotate JSON records for relevant cells, and Table-F1 calculated on double-annotated tables is used as the IAA score. A Table-F1 score of 91.0 underscores the reliability of the dataset.
DisCoMat (Gupta et al., 2022)
We experiment with DisCoMat, a dataset focusing on glass composition tables from Elsevier material science journals. The task is to extract tuples comprising (material, constituent, percentage, unit) from given tables. We adapt DisCoMat to fit our Schema-Driven IE framework by grounding the percentage element to numeric cells in the table and considering the other elements as attributes. The model is tasked to identify numeric cells representing constituent percentages and predict the associated three attributes. We refer readers to Gupta et al. (2022) for more details of DisCoMat. [^5] In the released corpus, tables are represented as matrices; we, therefore, transform them into CSV tables (using the pipe symbol ”—” as the delimiter) prior to feeding them into LLMs.
SWDE (Hao et al., 2011)
@@ 4 Experiments
We evaluate the capability of various LLMs to perform Schema-Driven IE, in addition to full fine-tuning using our benchmark. For ML and chemistry tables, we use a subset of 10 and 7 randomly sampled papers separately for model development, which facilitates the training of supervised baselines. For the two pre-existing datasets, we follow the data splits used in the original experiments.
4.1 Evaluation
To evaluate predicted JSON records we report F1 score against gold cell attributes that are exhaustively labeled by human annotators for each table. We report results both using exact match (EM), in addition to a threshold based on token-level similarity. The token similarity threshold is tuned on dev data to maximize alignment between our estimated model performance and performance measured using human judgments (see Appendix C for more details). We report F1 macro-averaged over tables, due to the wide variance in table sizes.
For DisCoMat and SWDE, we use similar metrics specified in their original papers to support comparison with prior work. We report Tuple-F1 (Gupta et al., 2022) for DisCoMat, where a predicted 4-element tuple is considered correct only if it exactly matches the gold tuple. For SWDE, we report Page-F1 (Hao et al., 2011), which measures the number of pages where the attributes are accurately predicted. [^6] Notably, SWDE primarily focuses on identifying textual HTML nodes containing attribute values rather than exact text spans, so we use token-level F1 to identify the most relevant HTML node for each extracted attribute.
To further validate our conclusions, we also present the results of full human evaluation of model outputs in §4.5.
@@ 4.2 Baselines & Implementation Details
We evaluate the capability of multiple LLMs to perform Schema-Driven IE, including API-based GPT-4 and GPT-3.5 models and open-source models, such as Llama3-8B-Instruct AI@Meta (2024), Llama2-Chat-13B (Touvron et al., 2023b), CodeLlama-instruct-13B (Rozière et al., 2023), StarCoder-15.5B (Li et al., 2023), LLaMA-7B (Touvron et al., 2023a), and Alpaca-7B (Taori et al., 2023). We also frame Schema-Driven IE as a TableQA problem, applying multi-choice and extractive QA prompts for template selection and cell attribute prediction, respectively. Furthermore, we also evaluate T5-11B (Raffel et al., 2020) and TaPas (Herzig et al., 2020), a table-specialized LM. For implementation details of InstrucTE and other methods, see Appendix D. [^7] We developed a rule-based method for chemistry tables based on the training set, which only achieved a Table-F1 score of 51.3, significantly lower than our proposed InstrucTE. Due to the substantial effort required to create specialized rule-based systems for each domain and the performance gap, we decided not to pursue this approach further.
@@ 4.3 Main Results
Despite a noticeable gap compared to API-based LLMs, open-source models show promising results across several domains. For example, CodeLlama-instruct-13B achieves 60.0 Table-F1 and 91.7 Page-F1 on ML tables and SWDE, respectively. Llama3-8B achieves 74.5 Table-F1 on chemistry tables.
@@ 4.4 Ablation Studies
Prompt Components & Error Recovery
@@ 4.5 Performance Analysis
To further verify our main conclusions from automatic evaluation and gain deeper insights into InstrucTE’s performance, we conduct a human evaluation and discuss a set of key questions.
Category % (#) Fine-grained Error Types Top 3 Affected Attributes
31.2 (48) gold answer in table, not predicted "Result:experimental_settings" (39.6%), "Result:training_data" (39.6%), "Result:model_settings" (20.8%)
False Negative 14.9 (23) gold answer in main text, not predicted "Result:training_data" (39.1%), "Hyper-parameter:model" (26.1%), "Hyper-parameter:dataset" (26.1%),
6.5 (10) gold answer predicted, wrong attribute "Result:experimental_settings" (100%)
19.5 (30) gold answer in main text, table header predicted "Result:training_data" (33.3%), "Result:task" (33.3%), "Result:model" (33.3%),
False Positive 11.7 (18) partial match, but misses important details "Result:test_data" (100%)
6.5 (10) gold answer in table caption, table header predicted "Result:metric" (100%)
6.5 (10) complete mismatch "Result:experimental_settings" (100%)
Propagated Errors 3.2 (5) select wrong record template "Other:type" (100%)
Table 3: Error analysis of InstrucTE (w/ GPT-4) for ML tables, inspecting 591 attribute predictions from 100 cell records sampled from 10 tables. For each fine-grained error type, we provide the error percentage, detailed error sources, and the top three affected attributes.
What errors are made by InstrucTE?
To understand where InstrucTE struggles, we conduct error analysis on GPT-4 predictions for ML tables. We sample 10 tables from the test set and 10 records for each table, comparing each attribute with the gold value. In total, we find 154 errors out of 591 attributes. We group errors into one of eight categories, listed in Table 3, with examples presented in Appendix F. For instance, one type of false positive error is when the gold attribute value is present in the table caption, but the model is distracted by a table header. Table 3 provides a detailed breakdown and includes the top three affected attributes within each error category. We find that the most common error occurs when the model fails to identify attributes present in the table (31.2%), particularly for experimental settings like 5-shot in Result records. Another major error is when attributes present in the accompanying text lead to either null predictions (14.9%) or incorrectly predicting a table header (19.5%). These errors highlight the challenges of Schema-Driven IE, where the model must understand nuances of table layouts and also effectively integrate information from surrounding text.
@@
Human Evaluation
Similar to our error analysis on ML tables, we manually inspect attribute predictions for 100 cell records for chemistry tables and DisCoMat, as well as 160 pages for SWDE, and report the prediction precision. Half of the sampled data are double-annotated, with the inter-annotator agreement score calculated as the F1 score between the two annotations. The statistics and results are provided in Table 9 in the appendix. The results show that InstrucTE achieves high precision across different datasets, ranging from 73.9 to 96.4, aligning with the performance under automatic metrics. Additionally, the high inter-annotator agreement scores (all above 90) indicate that the human evaluation is reliable and consistent.